OpenStudy (trwatkins1):

Subtract. –21 – (–13) A. –34 B. –8 C. 8 D. 34

OpenStudy (trwatkins1):

cal would this be the same?

OpenStudy (calculusxy):

Once again, the same thing applies: \[-21 + 13 \]

OpenStudy (michele_laino):

hint: \[\Large - 21 - \left( { - 13} \right) = - 21 + 13 = ...?\]

OpenStudy (trwatkins1):

oh ok i didnt see the front negitive

OpenStudy (anonymous):

Subtraction of a negative means to add a positive essentially. \[-21 - (-13) = is equivalent \to -21 + 13 = -8\]

OpenStudy (trwatkins1):

last time